### v2

# Bounding finite-image sequences of length $\omega^{k}$

Abstract.
Given a well-quasi-order $X$ and an ordinal $\alpha$, the set $s^{F}_{\alpha}(X)$ of transfinite sequences on $X$ with length less than $\alpha$ and with finite image is also a well-quasi-order, as proven by Nash-Williams. [11] Before Nash-Williams proved it for general $\alpha$, however, it was proven for $\alpha<\omega^{\omega}$ by Erdős and Rado. [5] In this paper, we revisit Erdős and Rado’s proof and improve upon it, using it to obtain upper bounds on the maximum linearization of $s^{F}_{\omega^{k}}(X)$ in terms of $k$ and $o(X)$, where $o(X)$ denotes the maximum linearization of $X$. We show that, for fixed $k$, $o(s^{F}_{\omega^{k}}(X))$ is bounded above by a function which can roughly be described as $(k+1)$-times exponential in $o(X)$. We also show that, for $k\leq 2$, this bound is not far from tight.
1. Introduction
Suppose $X$ is a well-quasi-order. Then there is a natural way of making $X^{*}$, the set of finite-length strings on $X$, into a well-quasi-order as well; this is Higman’s Lemma. [8] In 1965 C. Nash-Williams extended this from finite strings to transfinite sequences, with the qualification that one must restrict attention to sequences with finite image; we only allow a sequence if that sequence uses only finitely many symbols from $X$. [11] (Nash-Williams also showed that this restriction can be removed if $X$ is a better-quasi-order [12], but we will not consider this here.)
Given $X$ and an ordinal $\alpha$, we will define $s^{F}_{\alpha}(X)$ to be the set of finite-image sequences on $X$ with length less than $\alpha$. (So $s^{F}_{\omega}(X)$ is just $X^{*}$, since a string of finite length must certainly use only finitely many symbols.) The question then arises: Can we determine a nontrivial upper bound on the type of $s^{F}_{\alpha}(X)$, in terms of $\alpha$ and the type of $X$? (In the cases where we do not want a finiteness restriction, we will write $s_{\alpha}(X)$.)
Here by the type of a well-quasi-order $X$ we mean the largest order type of a linearization of $X$ (after quotienting out by equivalences); we denote it $o(X)$. This quantity was proven to exist by De Jongh and Parikh [4], who also showed that $o(X)$ can be characterized inductively as the smallest ordinal greater than $o(Y)$ for any proper lower set $Y$ of $X$. The theory has been rediscovered several times since then; indeed the term “type” comes from Kriz and Thomas. [9] This quantity is also known by other names, such as the maximum linearization.
Of course one could write down a trivial bound based on cardinality. But this is uninteresting. The question then is if one can do better; and, moreover, if one can come up with an upper bound that is at least reasonably tight. Schmidt claimed a nontrivial upper bound [14], but the proof has a hole which has never, to this author’s awareness, been repaired. [^1] The gap may be found on line 7 of page 34 of [14]; the second occurrence of $s_{\gamma}^{F}$ should simply be an $s_{\gamma}$, as the sequence need not have finite image, but without this finite-image condition, the proof cannot continue. Of course, Theorem 6, which is what’s being proven at that point, is true regardless; but the argument is later referred back to to prove Lemma 9 as part of the proof of the upper bound, and the argument is not valid, so Lemma 9 may not hold. Thanks to Andreas Weiermann for pointing this out to me. So this problem remains open.
Since the general problem seems difficult, we will restrict our attention to an easier special case, the case of small $\alpha$. It’s already been mentioned above that Higman’s Lemma, the case of $\alpha=\omega$, was proven before the full version of Nash-Williams’s theorem, and the problem of computing $o(X^{*})$ from $o(X)$ was solved by De Jong, Parikh, and Schmidt [4, 14]. But there’s more history to this problem inbetween these two endpoints. Before Nash-Williams’s theorem was proved, Erdős and Rado proved the special case of $\alpha<\omega^{\omega}$ [5]; before that, Rado proved the case of $\alpha=\omega^{3}$. [13] (A note, the result of Erdős and Rado is sometimes misstated as having been for $\alpha=\omega^{\omega}$, as in [10]; but in fact it was only for $\alpha<\omega^{\omega}$. A subsequent paper by Chopra and Pakhomov [2] will show how to obtain a bound on the type in the case $\alpha=\omega^{\omega}$.)
In this paper we will reexamine this proof of Erdős and Rado. This proof was written before the notion of $o(X)$ was ever defined, but we will see that we can extract from it an upper bound on $o(s^{F}_{\omega^{k}}(X))$. Actually, we will improve somewhat on this proof; the bound we will derive will be tighter than the one that would be obtained from a direct proof-mine.
Specifically:
Theorem 1.1.
For any fixed $k$ which is finite and nonzero, the type of $s^{F}_{\omega^{k}}(X)$ is bounded above by a function which is $(k+1)$-times exponential in $o(X)$ (in an appropriate sense to be made clear below). (See Theorem 3.14 for a precise version of this theorem.)
By contrast, a direct proof-mine of Erdős and Rado’s paper would result in a tower of $2k$ exponentials, rather than $k+1$.
Note, by the way, that this bound is much smaller than Schmidt’s claimed bound, so Schmidt’s claimed bound (which does not seem to have been intended to be tight in any way) is at least true in this case.
Unfortunately, it does not seem possible to extract a bound from Nash-Williams’s proof in the same way. So the problem of finding a bound for the general case, or even just for $o(s^{F}_{\omega^{\omega}}(X))$, remains open.
We will address the more general case of $\alpha<\omega^{\omega}$ in Section 3.1.
In Section 4, we will also examine the question of lower bounds; unfortunately, we are not able to show that our bound is reasonably tight in general, but we will at least show this when $n=2$ (the case of $n=1$ being already known due to the work of De Jongh, Parikh, and Schmidt [4, 14]).
More specifically:
Theorem 1.2.
There is a triply-exponential function $f(\beta)$ such that, for any $\beta$, there is some well-quasi-ordering $X$ with $o(X)=\beta$ and $o(s^{F}_{\omega^{2}}(X))\geq f(\beta)$.
Hopefully in the future this can be extended to larger $k$.
2. Preliminaries
We first recall some preliminaries about the behavior of $o(X)$. Much of the theory is often stated for well partial orders, with antisymmetry, rather than well-quasi-orders, without it, but of course this difference is immaterial; all statements should be taken as being up to equivalence in the relevant quasi-orders. That is to say, in a sense we are really talking about well partial orders, but we use well-quasi-orders here for convenience.
If $Y$ is a well-quasi-order and $X$ is a subset of $Y$, then $o(X)\leq o(Y)$, so more generally if $X$ embeds in $Y$, then $o(X)\leq o(Y)$. Moreover, if $X$ is a lower set in $Y$ and is also a proper subset of $Y$, then $o(X)<o(Y)$; indeed, $o(Y)$ is equal to the smallest ordinal greater than every $o(X)$, where $X$ ranges over all proper lower subsets of $Y$. [4] Relatedly, if there is a monotonic surjection from $Y$ onto $X$, then $o(X)\leq o(Y)$. (This is because $o(X)$ can also be characterized as the largest ordinal that $X$ has a monotonic surjection to; if $X$ has a monotonic surjection onto $Y$, then any ordinal that $Y$ surjects onto, $X$ will surject onto.) In particular, if $Y$ and $X$ are quasi-orders on the same set with the order $X$ being an extension of the order $Y$, then $o(X)\leq o(Y)$. [4]
Given two well-quasi-orders $X$ and $Y$, the disjoint union $X\amalg Y$ and Cartesian product $X\times Y$ are well-quasi-orders with $o(X\amalg Y)=o(X)\oplus o(Y)$ and $o(X\times Y)=o(X)\otimes o(Y)$, where $\oplus$ and $\otimes$ represent the natural (or “Hessenberg” [7, pp. 73–81]) addition and multiplication on the ordinals, distinct from their ordinary addition and multiplication. [4] These are also the addition and multiplication on the ordinals that one gets from the surreals. [3]
It’s also worth noting that $o(X)$ is a successor ordinal if and only if $X$ contains a maximal element (after equivalences); in this case one may remove that maximal element (and its equivalence class) to obtain $X^{\prime}$ with $o(X)=o(X^{\prime})+1$. [4]
Our central object of study here is $s^{F}_{\alpha}(X)$:
Definition 2.1.
Given a set $X$ and an ordinal $\alpha$, define $s_{\alpha}(X)$ to be the set of (possibly transfinite) sequences of length less than $\alpha$ taking values in $X$. (That is, functions $s:\beta\to X$, where $\beta$ is an ordinal less than $\alpha$.) Moreover, define $s^{F}_{\alpha}(X)$ to be the subset of $s_{\alpha}(X)$ consisting of those sequences with finite image.
In the case that $X$ is equipped with a well-quasi-order, one can use this to define a quasi-order on $s^{F}_{\alpha}(X)$, which Nash-Williams showed [11] is a well-quasi-order:
Definition 2.2.
Given a well-quasi-order $X$ and an ordinal $\alpha$, and given $s,t\in s^{F}_{\alpha}(X)$, define $s\leq t$ if there exists a strictly increasing function $\varphi:\operatorname*{dom}(s)\to\operatorname*{dom}(t)$ such that, for all $\alpha^{\prime}\in\operatorname*{dom}(s)$, one has $s(\alpha^{\prime})\leq t(\varphi(\alpha^{\prime}))$.
As noted earlier, $s^{F}_{\omega}(X)$ is just $X^{*}$, the set of strings on $X$. In this case, we have:
Theorem 2.3 (De Jongh, Parikh, Schmidt [4, 14]).
Let $X$ be a well-quasi-order. Then:
  • If $o(X)=0$, then $o(X^{*})=1$.
  • If $o(X)$ is nonzero and finite, then $o(X^{*})=\omega^{\omega^{o(X)-1}}$.
  • If $o(X)$ can be written as $\varepsilon+k$, where $\varepsilon$ is an epsilon number and $k$ is finite, then $o(X^{*})=\omega^{\omega^{o(X)+1}}$.
  • Otherwise, $o(X^{*})=\omega^{\omega^{o(X)}}$.
It’s worth observing here that in all of the constructions above – disjoint union, Cartesian product, strings on an alphabet – the type of the output depended on only on the type of the input. However this will not hold in general; in particular it will not hold for $s^{F}_{\alpha}(X)$ once $\alpha>\omega$. For instance, as observed by Schmidt [14], one has $o(s^{F}_{\omega+1}(2))=\omega^{\omega}2+1$ but $o(s^{F}_{\omega+1}(1\amalg 1))=\omega^{\omega}3+1$. Thus in general one can only put bounds on the result, rather than determine the result in terms of $o(X)$ alone.
Finally, given an well-quasi-order $X$, we’ll also consider a well-quasi-order on $\wp_{\mathrm{fin}}(X)$, the finite power set of $X$. There are actually multiple such quasi-orders in common use, but we will only concern ourselves with one of them.
Definition 2.4.
Given $S,T\in\wp_{\mathrm{fin}}(X)$, we define $S\leq_{\mathrm{m}}T$ if for each $s\in S$ there is some $t\in T$ such that $s\leq t$.
Notation 2.5.
Since we are only using $\leq_{\mathrm{m}}$, and are not using any of the other orders on $\wp_{\mathrm{fin}}(X)$, we will simply write $S\leq T$ rather than $S\leq_{\mathrm{m}}T$.
This is also a well-quasi-order, as follows from Higman’s Lemma. Once again we can put a bound on its type:
Theorem 2.6 (Abriola et. al. [1]).
If $X$ is a well-quasi-order, then
$o(\wp_{\mathrm{fin}}(X))\leq 2^{o(X)}.$
Moreover, this bound is tight; given an ordinal $\alpha$, there exists a well partial order $H_{\alpha}$ with $o(H_{\alpha})=\alpha$ and $o(\wp_{\mathrm{fin}}(H_{\alpha}))=2^{\alpha}$.
For convenience, we will also make the following definition:
Definition 2.7.
If $X$ is a set (particularly, a well-quasi-order), we define
$\wp^{\prime}_{\mathrm{fin}}(X):=\wp_{\mathrm{fin}}(X)\setminus\{\emptyset\}.$
Observe that $o(\wp^{\prime}_{\mathrm{fin}}(X))=-1+o(\wp_{\mathrm{fin}}(X))$; so, they are equal if $X$ is infinite, and $o(\wp^{\prime}_{\mathrm{fin}}(X))$ is one smaller if $X$ is finite.
3. Proof of the upper bound
In this section we prove the upper bound. As mentioned above, the proof is essentially similar to Erdős and Rado’s earlier proof [5] that did not provide an upper bound. The proof proceeds by induction on $k$, starting from the already-known case of $k=1$.
However, we make two alterations to Erdős and Rado’s approach. The first is is in how we approach sequences of length $\omega^{k}$; Erdős and Rado’s proof decomposed these by considering them as a sequence of length $\omega^{k-1}$ over an alphabet consisting of sequences of length $\omega$. We reverse this, and consider them as a sequence of length $\omega$ over an alphabet consisting of sequences of length $\omega^{k-1}$. I do not believe this change makes any substantial difference; I simply found the proof easier to carry out in this form.
The second change, however, is what allows us to derive a better bound than the one implicit in their proof. Where Erdős and Rado simply iterated the operation $X\mapsto X^{*}$, resulting in a tower of $2n$ exponentials, we will use this operation only once, and rely on $\wp_{\mathrm{fin}}$ (or rather $\wp^{\prime}_{\mathrm{fin}}$) for the rest, resulting in a tower of only $n+1$ exponentials.
We start by examining, what do sequences of length $\omega$ look like? This question was answered by Erdős and Rado, but we will repeat the proofs here. Let us make a definition:
Definition 3.1.
If $X$ is a well-quasi-order, let $e^{F}_{\alpha}(X)$ be the subset of $s^{F}_{\alpha+1}(X)$ consisting of sequences of length exactly $\alpha$. (Similarly, $e_{\alpha}(X)$ will be the subset of $s_{\alpha+1}(X)$ consisting of sequences of length exactly $\alpha$.)
It’s also useful here to know about indecomposable sequences:
Definition 3.2.
A sequence over a well-quasi-order $X$ is said to be indecomposable [6, Section 7.6.1] if it is equivalent to all of its nonempty tails. Note that the length of an indecomposable sequence must be an indecomposable ordinal, that is, a power of $\omega$. We will denote the set of indecomposable finite-image sequences of length exactly $\alpha$ over an alphabet $X$ by $z^{F}_{\alpha}(X)$.
Note that if a sequence is indecomposable, then so is any other sequence equivalent to it, since if $s\equiv t$ and $s$ indecomposable, then given a tail $t^{\prime}$ of $t$, using $s\leq t$ one may find a tail $s^{\prime}$ of $s$ with $s^{\prime}\leq t^{\prime}$, so $t\leq s\leq s^{\prime}\leq t^{\prime}$.
Proposition 3.3 (Erdős, Rado [5]).
Let $X$ be a well-quasi-order. Define the map $\varphi_{X}:\wp^{\prime}_{\mathrm{fin}}(X)\to z^{F}_{\omega}(X)$ as follows: To take $\varphi_{X}(S)$, write the elements of $S$ in an arbitrary order, so $S=\{s_{0},\ldots,s_{p-1}\}$. Then define $\varphi_{X}(S)=(s_{0}s_{1}\cdots s_{p-1})^{\omega}$. Then $\varphi_{X}$ is well-defined up to equivalences and monotonic.
Proof.
First, observe that the outputs are all indecomposable, since any nonempty tail of $(s_{0}s_{1}\cdots s_{p-1})^{\omega}$ is of the form $s_{r}\cdots s_{p-1}(s_{0}s_{1}\cdots s_{p-1})^{\omega}$ for some $r$ and so contains $(s_{0}s_{1}\cdots s_{p-1})^{\omega}$ in addition to being contained in it.
Now, we show monotonicity and well-definedness. Suppose $S=\{s_{0},\ldots,s_{p-1}\}$ and $T=\{t_{0},\ldots,t_{q-1}\}$, and suppose that for each $0\leq i<p$, there is some $j(i)$ such that $s_{i}\leq t_{j(i)}$. We wish to show that
$(s_{0}s_{1}\cdots s_{p-1})^{\omega}\leq(t_{0}t_{1}\cdots t_{q-1})^{\omega}.$
Write
$(t_{0}t_{1}\cdots t_{q-1})^{\omega}=((t_{0}t_{1}\cdots t_{q-1})^{p})^{\omega}.$
Certainly
$s_{0}s_{1}\cdots s_{p-1}\leq(t_{0}t_{1}\cdots t_{q-1})^{p},$
since each $s_{i}$ is less than or equal to the $i$’th copy of $t_{j(i)}$. Extending this implies
$(s_{0}s_{1}\cdots s_{p-1})^{\omega}\leq((t_{0}t_{1}\cdots t_{q-1})^{p})^{\omega}=(t_{0}t_{1}\cdots t_{q-1})^{\omega},$
as required.
Applying this with $S=T$ then shows that $\varphi_{X}$ is well-defined up to equivalences, and applying it for arbitrary $S$ and $T$ shows that $\varphi_{X}$ is monotonic. ∎
Proposition 3.4 (Erdős, Rado [5]).
Let $X$ be a well-quasi-order. Define the map $\varphi_{X}:X^{*}\times\wp^{\prime}_{\mathrm{fin}}(X)\to e^{F}_{\omega}(X)$ by $\varphi_{X}(t,S)=t\varphi_{X}(S)$. Then this $\varphi_{X}$ is well-defined up to equivalences; is monotonic; and is surjective up to equivalences. Moreover, given $s\in e^{F}_{\omega}(X)$, we may take $S$ to consist of all elements of $X$ that occur infinitely often in $s$.
You will notice here the reuse of the symbol $\varphi_{X}$; this is deliberate. We will soon be defining many maps of the form $\varphi_{X}:F(X)\to s^{F}_{\alpha}(X)$, each for a different $F(X)$. We could denote these $\varphi_{X,F(X)}$, but since there will be no ambiguity, we simply write $\varphi_{X}$ for all of them, which will make things easier.
Proof.
Since $\varphi_{X}:\wp^{\prime}_{\mathrm{fin}}(X)\to s^{F}_{\omega}(X)$ is well-defined up to equivalences and monotonic, it follows immediately that $\varphi_{X}:X^{*}\times\wp^{\prime}_{\mathrm{fin}}(X)\to e^{F}_{\omega}(X)$ is as well. The problem, then, is to show that it is surjective.
Suppose $s=s_{0}s_{1}\cdots\in e^{F}_{\omega}(X)$. Let $S=\{s_{0},s_{1},\ldots\}$; then $S$ is a finite set. Let $T\subseteq S$ be the set of elements of $S$ that occur infinitely often in $S$. Because $S\setminus T$ is finite, and each element of it occurs only finitely often, there must be some index $r$ such that, for all $i\geq r$, $s_{i}\in T$. Let $s^{\prime}=s_{r}s_{r+1}\cdots$.
We want to show, then, that $\varphi_{X}(s_{0}\cdots s_{r-1},T)$ is equivalent to $s$. Since the initial $r$ elements are equal, it suffices to show that $s^{\prime}$ is quivalent to $\varphi_{X}(T)$.
Say $T=\{t_{0},\ldots,t_{p-1}\}$. Then for each $i\geq r$ we have $s_{i}\in T$ and therefore $s_{i}\leq t_{0}\cdots t_{p-1}$; this then implies
$s^{\prime}\leq(t_{0}\cdots t_{p-1})^{\omega}=\varphi_{X}(T).$
For the reverse, to show $\varphi_{X}(T)\leq s^{\prime}$, we can inductively define a strictly increasing $j:\omega\to\omega$ such that
$(t_{0}\cdots t_{p-1})^{\omega}_{i}\leq s^{\prime}_{j(i)}.$
Assume that we have defined $j(i^{\prime})$ for all $i^{\prime}<i$, and we want to define $j(i)$. Now, $\varphi_{X}(T)_{i}$ is some element of $T$, which means it occurs infinitely often in $s^{\prime}$. Therefore, it occurs at least once with index greater than $j(i^{\prime})$ for any $i^{\prime}<i$. So, define $j(i)$ to be such an index.
Then
$(t_{0}\cdots t_{p-1})^{\omega}_{i}\leq s^{\prime}_{j(i)}$
as required, showing that $\varphi_{X}(T)\leq s^{\prime}$. So the two are equivalent, and $\varphi_{X}$ is surjective up to equivalences. Moreover, $T$ was selected in the manner required. ∎
Corollary 3.5.
Let $X$ be a well-quasi-order. Then $\varphi_{X}:\wp^{\prime}_{\mathrm{fin}}(X)\to z^{F}_{\omega}(X)$ is surjective up to equivalences. Moreover, a length-$\omega$ sequence is indecomposable if and only if each of its elements occur infinitely often, and if and only if it is equivalent it $\varphi_{X}(S)$, where $S$ is its image.
Proof.
By Proposition 3.4, if $s\in z^{F}_{\omega}(X)$, then $s=s^{\prime}\varphi_{X}(S)$ for some $s^{\prime}\in X^{*}$ and $S\in\wp^{\prime}_{\mathrm{fin}}(X)$. But this means that $\varphi_{X}(S)$ is a nonempty tail of $s$, and $s$ was assumed indecomposable, proving the surjectivity claim.
Moreover, if each element in a sequence of length $\omega$ occurs infinitely often, then by Proposition 3.4 it is equivalent to $\varphi_{X}(S)$, where $S$ is its image, and thus by Proposition 3.3 is indecomposable; and if it is indecomposable, then each element occurs infinitely often by the previous paragraph, so all of these conditions are equivalent. ∎
So now we know what sequences of length $\omega$ look like; in order to get a handle on sequences of length $\omega^{k}$, however, we’re going to have to make the domain of our mapping a bit bigger.
Definition 3.6.
Let $X$ be a well-quasi-order. Define $P_{k}(X)$ and $Q_{k}(X)$ inductively as follows:
  1. (1)
    $P_{0}(X)=X$
  2. (2)
    $Q_{k}(X)=\coprod_{i=0}^{k-1}P_{i}(X)$
  3. (3)
    For $k>0$, $P_{k}(X)=\wp^{\prime}_{\mathrm{fin}}(Q_{k}(X))$
Note that all of these are well-quasi-orders, and their types may be bounded (in terms of $o(X)$) by application of Theorem 2.6. Note also that $Q_{0}(X)$ is the empty set.
Now, just as we mapped $\wp^{\prime}_{\mathrm{fin}}(X)$ and $X^{*}\times\wp^{\prime}_{\mathrm{fin}}(X)$ to $e^{F}_{\omega}(X)$, we’ll do the same with these to cover $s^{F}_{\omega^{k}}(X)$ and $e^{F}_{\omega^{k}}(X)$.
Definition 3.7.
Let $X$ be a well-quasi-order. Define a map $\varphi_{X}:P_{k}(X)\to s^{F}_{\omega^{k}+1}(X)$, and (for $k>0$) a map $\varphi_{X}:Q_{k}(X)\to s^{F}_{\omega^{k-1}+1}(X)$, as follows. (Again, we’ll continue using the same symbol for all of these.)
  1. (1)
    $\varphi_{X}:P_{0}(X)\to s^{F}_{2}(X)$ will simply be the inclusion map $X\to X\cup\{\varepsilon\}$.
  2. (2)
    $\varphi_{X}:Q_{k}(X)\to s^{F}_{\omega^{k-1}+1}(X)$ will simply be the disjoint union of $\varphi_{X}$ on each of the $P_{k}(X)$ that make it up.
  3. (3)
    To define $\varphi_{X}:P_{k}(X)\to s^{F}_{\omega^{k}+1}(X)$ for $k>0$, suppose $S\in\wp^{\prime}_{\mathrm{fin}}(Q_{k}(X))=\{s_{0},s_{1},\ldots,s_{p-1}\}$; then
    $\varphi_{X}(S)=(\varphi_{X}(s_{0})\cdots\varphi_{X}(s_{p-1}))^{\omega}.$
Note how case (3) in Definitions 3.6 and 3.7 is connected to Propositions 3.3 and 3.4; we take finite sets to express sequences of the form $s^{\omega}$. Thus, for instance, an element $x\in X$ is quite different from $\{x\}\in\wp^{\prime}_{\mathrm{fin}}(X)$, as $\phi_{X}(x)=x$, while $\phi_{X}(\{x\})=x^{\omega}$.
Proposition 3.8.
The map $\varphi_{X}:P_{k}(X)\to s^{F}_{\omega^{k}+1}(X)$ and, for $k>0$, the map $\varphi_{X}:Q_{k}(X)\to s^{F}_{\omega^{k-1}+1}(X)$ are well-defined up to equivalence and are monotonic. Also, the image of the each map consists of indecomposable sequences.
Proof.
We induct on $k$, alternating between $P$ and $Q$ in the order $P_{0}(X)$, $Q_{1}(X)$, $P_{1}(X)$, $Q_{2}(X)$, etc. Obviously for each $k$, if it is true for $P_{\ell}(X)$ for all $\ell<k$ then it is true for $Q_{k}(X)$. And clearly it is true for $P_{0}(X)$. That leaves the case of $P_{k}(X)$ with $k>0$.
So for $P_{k}(X)$ with $k>0$, assume the statement is true for $Q_{k}(X)$. By Proposition 3.3, the map
$\varphi_{Q_{k}(X)}:\wp^{\prime}_{\mathrm{fin}}(Q_{k}(X))\to z^{F}_{\omega}(Q_{k}(X))$
is monotonic and well-defined up to equivalences. That is, given $S,T\in P_{k}(X)$, we have $\varphi_{Q_{k}(X)}(S)\leq\varphi_{Q_{k}(X)}(T)$, where here we are comparing length-$\omega$ sequences over $Q_{k}(X)$, using the order on $Q_{k-1}(X)$.
Now, if we take $\varphi_{Q_{k}(X)}(S)$ and apply to each of its elements the map $\varphi_{X}$, and concatenate all of these in place, we obtain $\varphi_{X}(S)$. But since $\varphi_{X}:Q_{k}(X)\to s^{F}_{\omega^{k-1}+1}(X)$ is monotonic, and concatenation (including infinitary concatenation) is monotonic, that means that overall we have described a monotonic operation, and $\varphi_{X}:P_{k}(X)\to s^{F}_{\omega^{k}+1}(X)$ is monotonic.
Finally, we most show indecomposability. It suffices to handle $P_{k}(X)$, as the image of $Q_{k}(X)$ is a union of images of $P_{i}(X)$. If $k=0$, then the image of $P_{0}(X)=X$ consists of sequences of length $1$, which are necessarily indecomposable. While if $k>0$, then each sequence in the image of $P_{k}(X)$ is of the form $s^{\omega}$ for some sequence $s$, and is therefore also indecomposable. ∎
Definition 3.9.
Let $X$ be a well-quasi-order. We define $\varphi_{X}:Q_{k}(X)^{*}\to s^{F}_{\omega^{k}}(X)$ by $\varphi_{X}(a_{1}\cdots a_{n})=\varphi_{X}(a_{1})\cdots\varphi_{X}(a_{n})$.
Definition 3.10.
Let $X$ be a well-quasi-order. We define $\varphi_{X}:Q_{k}(X)^{*}\times P_{k}(X)\to s^{F}_{\omega^{k}+1}(X)$ by $\varphi_{X}(a,b)=\varphi_{X}(a)\varphi_{X}(b)$.
Theorem 3.11.
The maps $\varphi_{X}:Q_{k}(X)^{*}\to s^{F}_{\omega^{k}}(X)$, $\varphi_{X}:Q_{k}(X)^{*}\times P_{k}(X)\to s^{F}_{\omega^{k}+1}(X)$ are well-defined up to equivalence and are monotonic. In addition, the former map is surjective up to equivalences, and the image
$\varphi_{X}(Q_{k}(X)^{*}\times P_{k}(X))$
contains $e^{F}_{\omega^{k}}(X)$ up to equivalences. Finally, the image $\varphi_{X}(P_{k}(X))$ contains $z^{F}_{\omega^{k}}(X)$ up to equivalences.
Note that in the case $k=1$, this is covered by Propositions 3.3 and 3.4.
Proof.
That these maps are monotonic and well-defined up to equivalences follows from Proposition 3.8. The problem, then, is to prove the statements about their images.
In fact, it suffices to prove these statements under the assumption that $X$ is finite. To see this, observe that since any $s$ in either $s^{F}_{\omega^{k}}(X)$ or $e^{F}_{\omega^{k}}(X)$ uses only finitely many elements from $X$, then if we let $X^{\prime}\subseteq X$ be this finite image of $s$, to prove the statement for $s$ it suffices to prove the statement with $X^{\prime}$ in place of $X$. So, if we prove it for all finite $X^{\prime}$ and $s$, we prove it for all $s$, and therefore for all $X$ and $s$, i.e., we have proved it for all $X$ regardless of finiteness. As such, for the rest of this proof, we assume $X$ is finite.
We will induct on $k$; assume the statements are true for all $\ell<k$. Then for $Q_{k}(X)^{*}$, say we are given $s\in s^{F}_{\omega^{k}}(X)$. If $s=\varepsilon$, then $s=\varphi_{X}(\varepsilon)$, so assume $s\neq\varepsilon$. Let $\alpha=\operatorname*{dom}(s)<\omega^{k}$, and write $\alpha$ in Cantor normal form as $\alpha=\omega^{\ell_{0}}+\ldots+\omega^{\ell_{r}}$, where $\ell_{0}\geq\ldots\geq\ell_{r}$. Then we may correspondingly split up $s$ as $s_{0}\cdots s_{r}$ with $\operatorname*{dom}(s_{i})=\omega^{\ell_{i}}$. Since each $\ell_{i}<k$, we may apply the inductive hypothesis to conclude that each $s_{i}$ is (up to equivalences) in the image of the map
$\varphi_{X}:Q_{\ell_{i}}(X)^{*}\times P_{\ell_{i}}(X)\to s^{F}_{\omega^{\ell_{i}}+1}(X).$
If we write $s_{i}=\varphi_{X}(t_{i},t^{\prime}_{i})$, then we may define
$t=t_{0}t^{\prime}_{0}\cdots t_{r}t^{\prime}_{r};$
since, for $\ell<k$, we have $Q_{\ell}(X)\subseteq Q_{k}(X)$ and $P_{\ell}(X)\subseteq Q_{k}(X)$, we have $t\in Q_{k}(X)^{*}$. Moreover, $\varphi_{X}(t)=s$, proving the statement for $Q_{k}(X)^{*}$.
For $Q_{k}(X)^{*}\times P_{k}(X)$, say now we are given $s\in e^{F}_{\omega^{k}}(X)$. If $k=0$, then $s$ consists of just a single element $x\in X$, and so we can write $s=\varphi_{X}(\varepsilon,x)$.
If $k>0$, then we may decompose $s$ as $s=s_{0}s_{1}s_{2}\cdots$, where $\operatorname*{dom}(s_{i})=\omega^{k-1}$. We now apply the inductive hypothesis to write $s_{i}=\varphi_{X}(w_{i},t_{i})$ (up to equivalences) where $w_{i}\in Q_{k-1}(X)^{*}$ and $t_{i}\in P_{k-1}(X)$. For convenience, define
$Y=\wp^{\prime}_{\mathrm{fin}}(Q_{k-1}(X))\times P_{k-1}(X)=(P_{k-1}(X))^{2}.$
So consider the sequence
$v:=(\operatorname*{im}(w_{i}),t_{i})_{i<\omega}$
of length $\omega$ over the alphabet $Y$. We wish to apply Proposition 3.4 to $v$, but first we need to know that $t$ uses only finitely many elements of $Y$. This is where we apply our assumption that $X$ is finite; since $X$ is finite, it is easily seen that $Y$ is also finite. Therefore, there is no question that $v$ uses only finitely many elements of $Y$.
With that settled, we may now apply Proposition 3.4 to $v$ to obtain $u\in Y^{*}$ and $S\in\wp^{\prime}_{\mathrm{fin}}(Y)$ such that $\varphi_{Y}(u,S)=v$ (up to equivalences).
This means that, for some $p$,
$u=(\operatorname*{im}(w_{0}),t_{0}),(\operatorname*{im}(w_{1}),t_{1}),\ldots,(\operatorname*{im}(w_{p-1}),t_{p-1}).$
Recall here that $w_{i}\in Q_{k-1}(X)^{*}$ and $t_{i}\in P_{k-1}(X)$, and
$Q_{k-1}(X)\cup P_{k-1}(X)\subseteq Q_{k}(X),$
so if we write
$u^{\prime}=w_{0}t_{0}\cdots w_{p-1}t_{p-1},$
then $u^{\prime}\in Q_{k}(X)^{*}$. Note that $\varphi_{X}(u^{\prime})=s_{0}\cdots s_{p-1}$.
Now we consider $S$, which we will write as $S=\{(T_{0},S_{0}),\ldots,(T_{q-1},S_{q-1})\}$. Let $T$ be the union of the $T_{i}$, and let
$U=T\cup\{S_{0},\ldots,S_{q-1}\}\in\wp^{\prime}_{\mathrm{fin}}(Q_{k}(X)).$
For convenience once again, define $Z=Q_{k}(X)$.
Let $s^{\prime}=s_{p}s_{p+1}\cdots$, and let
$t=w_{p}\varphi_{Z}(t_{p})w_{p+1}\varphi_{Z}(t_{p+1})\cdots\in Z^{*};$
we wish to show that, up to equivalences, $\varphi_{Z}(U)=t$. From above, we know that
$t_{p}t_{p+1}\cdots=\varphi_{Z}(\{S_{0},\ldots,S_{q-1}\})=(S_{0}S_{1}\cdots S_{q-1})^{\omega}$
and that the images of the $w_{i}$ also (up to equivalence) form a repeating sequence $T_{0},T_{1},\ldots,T_{q-1},T_{0},T_{1}$, …; recall here that each such image is an element of the set $\wp^{\prime}_{\mathrm{fin}}(Q_{k-1}(X))\subseteq\wp^{\prime}_{\mathrm{fin}}(Z)$.
Putting this together, we see that $t$ (a length-$\omega$ sequence over $Z$) has image equal to $U$, and that each element in it occurs infinitely often; each $S_{i}$ occurs infinitely often as one of the $t_{i}$, each element of $T$ occurs infinitely often among the $w_{i}$ (since it occurs in some $\operatorname*{im}(w_{i})$, and there are infinitely many $w_{i}$ with that same image), and these are all the elements that occur. Therefore, by Corollary 3.5, $t$ is (as a sequence over $Z$) indecomposable and equivalent to $\varphi_{Z}(U)$.
So if we now pass from $Z$ down to $X$, we conclude by the monotonicity of concatenation that $s^{\prime}$, which recall is given by
$s^{\prime}=s_{p}s_{p+1}\cdots=\varphi_{X}(w_{p})\varphi_{X}(t_{p})\varphi_{X}(w_{p+1})\varphi_{X}(t_{p+1})\cdots,$
is equivalent to $\varphi_{X}(U)$; and since we have shown that $\varphi_{X}(u^{\prime})=s_{0}\cdots s_{p-1}$, we conclude that $\varphi_{X}(u^{\prime},U)=s$. Since $u^{\prime}\in Q_{k}(X)^{*}$ and $U\in P_{k}(X)$, that proves the required surjectivity in this case.
Finally, we handle the case of indecomposable sequences. Say we are given $s\in z^{F}_{\omega^{k}}(X)$. By the preceding argument, one may write (up to equivalences) $s=\phi_{X}(u^{\prime})\phi_{X}(S)$ for some $(u^{\prime},S)\in Q_{k}(X)^{*}\times P_{k}(X)$. But by assumption, $s$ is indecomposable, and $\phi_{X}(S)$ is a nonempty tail of it; so $s$ is equivalent to $\phi_{X}(S)$, concluding the proof of the theorem. ∎
Remark 3.12.
Based on Proposition 3.4, which shows that every length-$\omega$ finite-image sequence has an indecomposable tail, one might wonder whether the same is true for more general finite-image sequences; in fact, this is true (see [6, Section 7.6.4]), but we will not show it here.
We can then write out explicitly the bound this gets us.
Definitions 3.13.
Define sequences of functions $p_{k}$ and $q_{k}$ recursively via:
  1. (1)
    $p_{0}(\beta)=\beta$
  2. (2)
    $q_{k}(\beta)=\bigoplus_{i=0}^{k-1}p_{i}(\beta)$
  3. (3)
    For $k>0$, $p_{k}(\beta)=-1+2^{q_{k}(\beta)}$.
Theorem 3.14.
Let $h$ be the function such that $o(X^{*})=h(o(X))$ (see Theorem 2.3). Then, for any well-quasi-order $X$ and any whole number $k$, one has
$o(s^{F}_{\omega^{k}}(X))\leq h(q_{k}(o(X))),$
$o(e^{F}_{\omega^{k}}(X))\leq h(q_{k}(o(X)))\otimes p_{k}(o(X)),$
and
$o(z^{F}_{\omega^{k}}(X))\leq p_{k}(o(X)).$
This is the more explicit version of Theorem 1.1.
Proof.
This follows immediately from Theorem 3.11 together with Theorem 2.3 as well as the other preliminaries mentioned in Section 2. ∎
Example 3.15.
Let’s examine what this gets us for $s^{F}_{\omega^{k}}(x)$ for small $k$. For the trivial case of $k=0$, we get $o(s^{F}_{1}(X))\leq 1$, as we should, and for the well-known case of $k=1$, we get
$o(s^{F}_{\omega}(X))\leq h(o(X))\leq\omega^{\omega^{o(X)+1}},$
again as we should.
Meanwhile, for our first new case, $k=2$, we get
$o(s^{F}_{\omega^{2}}(X))\leq h((-1+2^{o(X)})\oplus o(X))\leq\omega^{\omega^{(-1+2^{o(X)})\oplus o(X)+1}}.$
Let’s now consider what happens if $o(X)=2$ (that is, $|X|=2$). In this case, the bound becomes
$o(s^{F}_{\omega^{2}}(X))\leq h(5)=\omega^{\omega^{4}}.$
Now, if $|X|=2$, then either $X$ is a chain of size $2$, or an antichain of size $2$. In the case of a chain, say $X=\{0,1\}$, this is certainly an overestimate; the actual type would be at most $h(4)=\omega^{\omega^{3}}$, as $(01)^{\omega}$ and $1^{\omega}$ are equivalent, so one has only $4$ distinct building blocks, not $5$. In the case of an antichain, by contrast, it is possible that this bound is tight, although I have not verified whether it is.
3.1. The more general case of $\alpha<\omega^{\omega}$
Let us briefly address the more general case of $s^{F}_{\alpha}(X)$ and $e^{F}_{\alpha}(X)$ for $\alpha<\omega$. Of course, if one only wants crude bounds, one can simply find some $k$ such that $\alpha\leq\omega^{k}$ and then apply Theorem 3.14, but it’s not hard to do better, so let’s address that.
For convenience, let’s make some definitions first.
Definitions 3.16.
As above, we’$l$ define $h$ to be the function such that $o(X^{*})=h(o(X))$ (see Theorem 2.3). We’ll also define
$f_{k}(\beta)=h(q_{k}(\beta))$
and
$g_{k}(\beta)=h(q_{k}(\beta))\otimes p_{k}(\beta),$
so that (per Theorem 3.14) we have $o(s^{F}_{\omega^{k}}(X))\leq f_{k}(o(X))$ and $o(e^{F}_{\omega^{k}}(X))\leq g_{k}(o(X))$.
With this, the case of $e^{F}_{\alpha}(X)$ is quite simple:
Theorem 3.17.
Suppose $\alpha=\omega^{k_{0}}+\ldots+\omega^{k_{r}}$, where $\omega>k_{0}\geq\ldots\geq k_{r}$. Then
$o(e^{F}_{\alpha}(X))\leq\bigotimes_{i=0}^{r}g_{k_{i}}(o(X)).$
Proof.
Concatenation forms a monotonic surjection from $\prod_{i=0}^{r}e^{F}_{\omega^{k}}(X)$ to $e^{F}_{\alpha}(X)$. ∎
We can handle the case of $s^{F}_{\alpha}(X)$ with only slightly more complication.
Theorem 3.18.
Suppose $\alpha=\omega^{k_{0}}+\ldots+\omega^{k_{r}}$, where $\omega>k_{0}\geq\ldots\geq k_{r}$. Then
$o(s^{F}_{\alpha}(X))\leq\bigoplus_{i=0}^{r}\left(\left(\bigotimes_{j=0}^{i-1}g_{k_{j}}(o(X))\right)\otimes f_{k_{i}}(o(X))\right).$
Proof.
For any sequence of length less than $\alpha$, we can find the smallest $i$ so that its length is less than $\omega^{k_{0}}+\ldots+\omega^{k_{i}}$; then the sequence breaks down into sequences of length equal to $\omega^{k_{j}}$ (for $j<i$) and a sequence of length less than $\omega^{k_{i}}$. This gives us a monotonic surjection onto $s^{F}_{\alpha}(X)$ from
$\coprod_{i=0}^{r}\left(\left(\prod_{j=0}^{i-1}e^{F}_{\omega^{k_{j}}}(X)\right)\times s^{F}_{\omega^{k_{i}}}(X)\right).$
However, when there’s a nonzero finite part it’s actually possible to do a little better than this.
Definitions 3.19.
Define $P_{k}^{0}(X)=\wp_{\mathrm{fin}}(Q_{k}(X))$, i.e., $P_{k}^{0}(X)=P_{k}(X)\cup\{\emptyset\}$. Then extend $\varphi_{X}$ from $P_{k}(X)$ to $P_{k}^{0}(X)$ via $\varphi_{X}(\emptyset)=\varepsilon$. We similarly extend $\varphi_{X}$ from $Q_{k}(X)^{*}\times P_{k}(X)$ to $Q_{k}(X)^{*}\times P_{k}^{0}(X)$ in the obvious manner.
Proposition 3.20.
The map $\varphi_{X}:P_{k}^{0}(X)\to s^{F}_{\omega^{k}+1}(X)$ is well-defined up to equivalences. The map $\varphi_{X}:Q_{k}(X)^{*}\times P_{k}^{0}(X)\to s^{F}_{\omega^{k}+1}(X)$ is well-defined up to equivalences, monotonic, and surjective up to equivalences.
Proof.
Since we already know that $\varphi_{X}:P_{k}(X)\to s^{F}_{\omega^{k}+1}(X)$ is well-defined up to equivalences and monotonic by Proposition 3.8, the same obviously also holds on $P_{k}^{0}(X)$, which then implies it also holds (by Proposition 3.8 again) on $Q_{k}(X)^{*}\times P_{k}^{0}(X)$. So the only question is surjectivity.
Any element of $s^{F}_{\omega^{k}+1}(X)$ is either an element of $s^{F}_{\omega^{k}}(X)$ or $e^{F}_{\omega^{k}}(X)$. In the former case, by Theorem 3.11, it (up to equivalences) lies in the image of $Q_{k}(X)^{*}$, i.e., in the image of $Q_{k}(X)^{*}\times\{\emptyset\}$. In the latter case, again by Theorem 3.11, it (up to equivalences) lies in the image of $Q_{k}(X)^{*}\times P_{k}(X)$. Either way, up to equivalences, it lies in the image of $Q_{k}(X)^{*}\times P_{k}^{0}(X)$. ∎
Based on this, we define:
Definition 3.21.
Let $p^{+}_{k}(\beta)=1+p_{k}(\beta)$ and $g^{+}_{k}(\beta)=f_{k}(\beta)\otimes p^{+}_{k}(\beta)$.
And from this we conclude the following proposition and theorem:
Proposition 3.22.
If $X$ is a well-quasi-order, one has $o(s^{F}_{\omega^{k}+1}(X))\leq g^{+}_{k}(o(X))$.
Proof.
This follows immediately from Proposition 3.20. ∎
Theorem 3.23.
Suppose $\alpha=\omega^{k_{0}}+\ldots+\omega^{k_{r}}+\ell$, where $\omega>k_{0}\geq\ldots\geq k_{r}>0$ and $1\leq\ell<\omega$.
Then
$o(s^{F}_{\alpha}(X))\leq\bigoplus_{i=0}^{r-1}\left(\left(\bigotimes_{j=0}^{i-1}g_{k_{j}}(o(X))\right)\otimes f_{k_{i}}(o(X))\right)\oplus\\ \left(\left(\bigotimes_{j=0}^{r-1}g_{k_{j}}(o(X))\right)\otimes g^{+}_{k_{r}}(o(X))\right)\oplus\left(\left(\bigotimes_{j=0}^{r}g_{k_{j}}(o(X))\right)\otimes\left(\bigoplus_{t=1}^{\ell-1}o(X)^{\otimes t}\right)\right).$
Remark 3.24.
Note that the final term in this sum is not actually unique to this theorem; if one were to apply Theorem 3.18 in this case and expand it out, one would see it there too (since $f_{0}(\beta)=1$ and $g_{0}(\beta)=\beta)$). It is only the second-to-last term, that involves $g_{k_{r}}^{+}$, that is different.
Proof.
For any sequence of length less than $\alpha$, either its length is less than $\omega^{k_{0}}+\ldots+\omega^{k_{r-1}}$, or it is at least that long. In the former case, we apply Theorem 3.18. In the latter case, we can break it down into a sequence of length $\omega^{k_{0}}+\ldots+\omega^{k_{r-1}}$ (to which Theorem 3.17 applies) and a remainder of length less than $\omega^{k_{r}}+\ell$, i.e., at most $\omega^{k_{r}}+(\ell-1)$.
We can then break this down into the case where the length of the remainder is less than or equal to $\omega^{k_{r}}$, and the case where it is not, i.e., where it is equal to $\omega^{k_{r}}+t$ for some $1\leq t<\ell-1$.
Then the first case gives us the first part of the sum, the second case gives us the second part of the sum, and the third case gives us the third part of the sum. ∎
Example 3.25.
Let’s consider what this gets us when $\alpha=\omega^{2}+1$ and $o(X)=\omega$. If we apply Theorem 3.18, we see that
$o(s^{F}_{\omega^{2}+1}(X))\leq f_{2}(\omega)\oplus(f_{0}(\omega)\otimes g_{2}(\omega))=\omega^{\omega^{\omega 2}+2}+\omega^{\omega^{\omega 2}}.$
However, we apply Theorem 3.22, we see that in fact
$o(s^{F}_{\omega^{2}+1}(X))\leq g^{+}_{2}(\omega)=\omega^{\omega^{\omega 2}+2}.$
It’s probably possible to improve on this further, but we will stop here.
4. Lower bounds
We now turn to the question of lower bounds. We’ll prove our lower bounds on $s^{F}_{\omega^{2}}(X)$ by embedding another ordering into it.
Proposition 4.1.
Suppose $X$ is a well partial order with elements $v_{1},\ldots,v_{k}$ such that each $v_{i}$ is maximal in $X\setminus\{v_{i+1},\ldots,v_{k}\}$. Let $Y=X\setminus\{v_{1},\ldots,v_{k}\}$. Then there is an embedding of $(\wp^{\prime}_{\mathrm{fin}})^{k}(Y)$ into $e^{F}_{\omega^{k}}(X)$ (the image of which, for $k>0$, consists of indecomposable sequences) and an embedding of $(\wp^{\prime}_{\mathrm{fin}})^{k}(Y)^{*}$ into $s^{F}_{\omega^{k+1}}(X)$.
Remark 4.2.
This proposition can obviously be applied to e.g. $Y+k$ or $Y\amalg k$ or $Y$ together with antichain of size $k$, but this proposition can actually be applied to any well partial order $X$ with $o(X)=\beta+k$. Since (as mentioned in Section 2 and proved in [4]) any $X$ with $o(X)$ equal to a successor ordinal has a maximal element, if $o(X)=\beta+k$, then one can take a maximal element $v_{k}$, remove it, and iterate this process $k$ times to obtain $v_{1},\ldots,v_{k}$ as above.
Proof.
We’ll start by constructing the embedding
$\psi_{k}:(\wp^{\prime}_{\mathrm{fin}})^{k}(Y)\to z^{F}_{\omega^{k}}(Y\cup\{v_{1},\ldots,v_{k}\}).$
We construct this by induction on $k$. For $k=0$, $\psi_{0}:Y\to Y$ will simply be the identity map. Now, if we have defined $\psi_{\ell}$, and we have $S=\{T_{1},\ldots,T_{r}\}\in(\wp^{\prime}_{\mathrm{fin}})^{\ell+1}(Y)$, we define
$\psi_{\ell+1}(S)=(v_{\ell}\psi_{\ell}(T_{1})v_{\ell}\psi_{\ell}(T_{2})\cdots v_{\ell}\psi_{\ell}(T_{r}))^{\omega}.$
We wish to show that this is well-defined up to equivalences, and an embedding.
We show this by induction on $\ell$. By inductively applying Proposition 3.3, we can see that it’s well-defined up to equivalences and monotonic; it’s also easy to see that it’s indecomposable. And if $k=0$, it’s obviously also an embedding. Suppose it’s an embedding for $\ell$ and we wish to prove it for $\ell+1$. So suppose that $\psi_{\ell+1}(S)\leq\psi_{\ell+1}(S^{\prime})$. Since $v_{\ell}$ is maximal in $Y\cup\{v_{1},\ldots,v_{\ell}\}$, each $v_{\ell}$ that occurs in $\psi_{\ell+1}(S)$ must map to a $v_{\ell}$ in $\psi_{\ell+1}(S^{\prime})$. But since each $v_{\ell}\psi_{\ell}(T)$ (for $T\in(\wp^{\prime}_{\mathrm{fin}})^{\ell}(Y)$) begins with a $v_{\ell}$, this means that each $v_{\ell}\psi_{\ell}(T)$ (for $T\in S)$ must be wholly contained within a single $v_{\ell}\psi_{\ell}(T^{\prime})$ (for $T^{\prime}\in S^{\prime})$, and so $\psi_{\ell}(T)$ is wholly contained within $\psi_{\ell}(T^{\prime})$. Thus we conclude that for each $T\in S$, there is some $T^{\prime}\in S^{\prime}$ such that $\psi_{\ell}(T)\leq\psi_{\ell}(T^{\prime})$. Since $\psi_{\ell}$ is an embedding by the inductive hypothesis, we get that for each $T\in S$ there is some $T^{\prime}\in S^{\prime}$ such that $T\leq T^{\prime}$; this means that $S\leq S^{\prime}$ and so $\psi_{\ell+1}$ is an embedding as claimed.
We can now define $\psi^{\prime}_{k}:(\wp^{\prime}_{\mathrm{fin}})^{k}(Y)*\to s^{F}_{\omega^{k+1}}(Y\cup\{v_{1},\ldots,v_{k}\})$ by $\psi^{\prime}_{k}(S_{1}\cdots S_{r})=v_{k}\psi_{k}(S_{1})v_{k}\psi_{k}(S_{2})\cdots v_{k}\psi_{k}(S_{r})$. This is clearly monotonic, and, by a similar argument as to above, it is also an embedding. This proves the claim. ∎
Now, we’ll need a family of well partial orders $H_{\beta}$ such that $o(H_{\beta})=\beta$ and such that we get large enough types upon performing other operations on $H_{\beta}$. Fortunately, [1] provides us with one.
Definition 4.3.
Following Abriola et. al. [1], define the well partial orders $H_{\beta}$ recursively by
  1. (1)
    $H_{1}=1$
  2. (2)
    $H_{\omega}=\sum_{k<\omega}\coprod_{i<k}1$
  3. (3)
    For $\gamma>0$, $H_{\omega^{\omega^{\gamma}}}=\sum_{\delta<\omega^{\gamma}}H_{\omega^{\delta}}$
  4. (4)
    For $H_{\omega^{\gamma}}$ where $\gamma$ is not a power of $\omega$, write $\gamma=\omega^{\delta_{0}}+\ldots+\omega^{\delta_{r}}$ ($\delta_{i}$ weakly decreasing); then $H_{\omega^{\gamma}}=\prod_{i}H_{\omega^{\delta_{i}}}$, where the product is a lexicographic product.
  5. (5)
    If $\beta$ is not a power of $\omega$, write $\beta=\omega^{\gamma_{0}}+\ldots+\omega^{\gamma_{r}}$ ($\gamma_{i}$ weakly decreasing); then $H_{\beta}=\coprod_{i}H_{\omega^{\gamma_{i}}}$.
Here the sums are ordered sums, i.e., concatenation.
Remark 4.4.
There are any number of other similar ways one could construct a family $K_{\alpha}$ with the same properties as $H_{\alpha}$. For instance, one could define
  1. (1)
    $K_{1}=1$
  2. (2)
    $K_{\omega^{\gamma+1}}=\sum_{k<\omega}K_{\omega^{\gamma}k}$
  3. (3)
    $K_{\omega^{\gamma}}=\sum_{\gamma^{\prime}<\gamma}K_{\omega^{\gamma^{\prime}}}$ if $\gamma$ is a limit ordinal
  4. (4)
    If $\beta$ is not a power of $\omega$, write $\beta=\omega^{\gamma_{0}}+\ldots+\omega^{\gamma_{r}}$ ($\gamma_{i}$ weakly decreasing); then $K_{\beta}=\coprod_{i}K_{\omega^{\gamma_{i}}}$.
This is not always isomorphic, as can be seen by taking $\alpha=\omega^{\omega^{2}+\omega}$ ($H_{\omega^{\omega^{2}+\omega}}$ has two elements of height $\omega^{\omega^{2}}$, whereas $K_{\omega^{\omega^{2}+\omega}}$ has only one), but it would work just as well. We will stick to the family used by Abriola et. al., however.
Abriola et. al. showed:
Proposition 4.5 ([1, Proposition 4.2 and Theorem 4.3]).
For any $\beta$, one has $o(H_{\beta})=\beta$ and $o(\wp_{\mathrm{fin}}(H_{\beta}))=2^{\beta}$.
We’ll also need one other property of this family:
Proposition 4.6.
If $\beta\leq\beta^{\prime}$, then $H_{\beta}$ embeds in $H_{\beta}^{\prime}$.
Proof.
We induct on $\beta^{\prime}$. Write $\beta$ in Cantor normal form as $\gamma=\omega^{\gamma_{0}}+\ldots+\omega^{\gamma_{r}}$ with $\gamma_{i}$ weakly decreasing, and similarly write $\beta^{\prime}=\omega^{\gamma^{\prime}_{0}}+\ldots+\omega^{\gamma^{\prime}_{s}}$.
Since $\beta\leq\beta^{\prime}$, either $\gamma=\gamma^{\prime}_{i}$ for all $i$ and we have $r\leq s$, or there must be some smallest $k$ such that $\gamma_{k}<\gamma^{\prime}_{k}$. In the first case, the embedding is immediate by the construction.
In the second case, we can use $H_{\omega^{\gamma_{i}}}=H_{\omega^{\gamma^{\prime}_{i}}}$ for $i<k$; the problem then is to embed $\coprod_{i\geq k}H_{\omega^{\gamma_{i}}}$ inside $H_{\omega^{\gamma^{\prime}_{k}}}$.
Since we are in the case where $\gamma_{k}<\gamma^{\prime}_{k}$, we cannot have $\gamma^{\prime}_{k}=0$. If $\gamma^{\prime}_{k}$ is a power of $\omega$, then the embedding is immediate by the construction. If $\gamma^{\prime}_{k}=\delta+1$, then by the inductive hypothesis we may embed each $\gamma_{i}$ for $i>k$ into $H_{\delta}$; we can then embed the whole into $H_{\omega^{\gamma^{\prime}_{k}}}=H_{\omega^{\delta}}\cdot H_{\omega}$ since $H_{\omega}$ is infinite.
This leaves the case where $\gamma^{\prime}_{k}$ is a limit ordinal but not a power of $\omega$. In this case, we may choose some successor ordinal $\delta$ with $\gamma_{k}\leq\gamma^{\prime\prime}<\gamma^{\prime}_{k}$; then the whole embeds into $H_{\omega^{\gamma^{\prime\prime}}}$ by the above, and the problem is now to embed $H_{\omega^{\gamma^{\prime\prime}}}$ into $H_{\gamma^{\prime}_{k}}$.
So expand $\gamma^{\prime}_{k}$ into Cantor normal form as the sum of $\omega^{\delta^{\prime}_{i}}$, and $\gamma^{\prime\prime}$ into Cantor normal form as the sum of $\omega^{\delta^{\prime\prime}_{i}}$. By the product construction, we may once again cancel common terms at the top, meaning we may reduce to the case where $\gamma^{\prime}_{k}$ is a power of $\omega$. But this case was already handled; this completes the proof. ∎
We now define our “triply exponential” function that we will use as our lower bound.
Definition 4.7.
Define a function $u$ by $u(\beta)=h(-1+2^{\beta-1})$ if $\beta$ is a successor ordinal, $u(\beta)=\omega^{\omega^{2^{\beta}}}$ if $\beta$ is a limit ordinal, and $u(0)=1$.
We’ve defined this function so as to ensure it is continuous.
Proposition 4.8.
The function $u$ is continuous.
Proof.
On limit ordinals, $u(\beta)=\omega^{\omega^{2^{\beta}}}$, so it is continuous restricted to this subset; so it suffices to check that
$\lim_{k\to\omega}u(\beta+k)=u(\beta+\omega)$
when $\beta$ is a limit ordinal or 0. Since
$\omega^{\omega^{-2+2^{\beta+k-1}}}\leq u(\beta+k)\leq\omega^{\omega^{2^{\beta+k-1}+1}}$
when $k>0$ and both these bounds tend to $\omega^{\omega^{2^{\beta+\omega}}}$, it follows that the limit is $\omega^{\omega^{2^{\beta+\omega}}}$, as desired. ∎
Putting all this together yields the following theorem:
Theorem 4.9.
For any ordinal $\beta$, there exists a well partial order $X$ with $o(X)=\beta$ and $o(s^{F}_{\omega^{2}}(X))\geq u(\beta)$. Specifically, we may take $X=H_{\beta}$.
Proof.
If $\beta=0$, the claim is trivial. If $\beta$ is a successor ordinal, write $\beta=\beta^{\prime}+1$, so $H_{\beta}=H_{\beta}\amalg 1$. Then by Proposition 4.1, there is an embedding of $\wp^{\prime}_{\mathrm{fin}}(H_{\beta^{\prime}})^{*}$ into $s^{F}_{\omega^{2}}(H_{\beta})$. Then $o(s^{F}_{\omega^{2}}(H_{\beta}))\geq h(-1+2^{\beta^{\prime}})=u(\beta)$.
Finally, if $\beta$ is a limit ordinal, then we may assume an inductive hypothesis that the statement is true for all $\beta^{\prime}<\beta$. Since $H_{\beta^{\prime}}$ embeds in $H_{\beta}$ by Proposition 4.6, we get an embedding of $s^{F}_{\omega^{2}}(H_{\beta}^{\prime})$ into $s^{F}_{\omega^{2}}(H_{\beta})$. So
$o(s^{F}_{\omega^{2}}(H_{\beta}))\geq\sup_{\beta^{\prime}<\beta}o(s^{F}_{\omega^{2}}(H_{\beta^{\prime}}))\geq u(\beta^{\prime});$
since $u$ is continuous, this means
$o(s^{F}_{\omega^{2}}(H_{\beta}))\geq u(\beta).$
This proves the theorem. ∎
Finally, we also note:
Proposition 4.10.
Given an ordinal $\beta>0$, there exists a well partial order $X$ with $o(X)=\beta$ and $o(z^{F}_{\omega}(X))\geq-1+2^{\beta-1}$ if $\beta$ is a successor ordinal and $o(z^{F}_{\omega}(X))=-1+2^{\beta}$ if $\beta$ is a limit ordinal.
Proof.
This follows from Propositions 4.1 and 4.5 by similar reasoning to as above; if $\beta$ is a successor ordinal, one removes a maximal element and applies Proposition 4.1, while if it’s a limit ordinal, one takes limits. In the latter case we must also have equality as here the lower bound matches the upper bound from Theorem 3.14. ∎
We hope that Theorem 4.9 and Proposition 4.10 can be extended to the case of $\omega^{k}$ in the future. Note that the limiting step here is getting lower bounds on the type of iterations of $\wp_{\mathrm{fin}}$ beyond the first.
Acknowledgements
Thanks to A. D. Chopra and to an anonymous referee for various helpful comments.